If 8 = b^-2, what is the value of b?
Answer by Theo(13342)

your equation is 8 = b^-2.
since b^-2 = 1/b^2, this equation becomes 8 = 1/b^2.
multiply both sides of this equation by b^2 and divide both sides of this equation by 8 to get b^2 = 1/8.
take the square root of both sides of this equation to get b = plus or minus sqrt(1/8).

to confirm this solution is correct, replace b with plus or minus sqrt(1/8) and solve the equation.

put in your calculator (sqrt(1/8)^-2 and the calculator will tell you that the answer is 8.

put in your calculator (-sqrt(1/8)^-2 and the calculator will tell you that the answer is 8.

this confirms the solution is correct.

the solution is b = plus or minus sqrt(1/8).